OpenStudy (anonymous):

\[\int\limits_{0}^{2\Pi}\left| senx-cosx \right|dx\]

OpenStudy (anonymous):

Is this from 0 to 2pi?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Absolute value

OpenStudy (anonymous):

give me just a sec

OpenStudy (anonymous):

what senx is?

OpenStudy (anonymous):

sinx, sorry (I speak spanish, in spanish is seno , not sine :D)

OpenStudy (anonymous):

I'm getting 4squareroot of 2...I'll be right back...

OpenStudy (anonymous):

\[\int\limits\limits\limits_{0}^{\pi/4}+\int\limits\limits\limits_{\pi/4}^{\pi/2}+\int\limits\limits\limits_{\pi/2}^{3\pi/4}+\int\limits\limits\limits_{3\pi/4}^{2\pi}..\]

OpenStudy (anonymous):